OpenStudy (charlotte123):

Which proportion could be used to solve for the missing side length? a. s/5 = 22.5/10 b. 4/18 = 10/s c. 4/s = 10/18 d. s/5 = 10/22.5

OpenStudy (charlotte123):

Here is a picture:

Directrix (directrix):

,( d). s/5 = 10/22.5 --> Answer The triangles are similar and the lengths of corresponding sides are in proportion. So, 18/4 = 10/s = 22.5/5 From that extended proportion,( d). s/5 = 10/22.5 is equivalent to 10/s = 22.5/5.

Directrix (directrix):

Question?